40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 96 of 341  Not logged in ELOG logo
ID Date Author Type Category Subject
  12348   Thu Jul 28 16:43:01 2016 LydiaUpdateGeneralETMX aluminium standoff groove condition

 

I took some pictures with the digital microscope of the aluminum standoffs removed from ETMX. The first one had some leftover epoxy still attached, so I was able to distinguish which part of the groove was occupied by the wire. A better microscope would help (this one has a maximum magification of 80, 200 or so would be much better) but I was still able to see what looks like a second minimum inside the groove at the wire location (see Attachments 1 and 2). The bottom edge of the standoff shows the profile of the groove on the opposite side from the glue. I took several photos with different lighting angles and at different locations on the microscope stage and convinced myself that this was not just an artificial effect. I also took photos of the groove in a different place and did not see this feature (Attachment 3).

The other standoff in the same container had no visible damage to the groove or to the body of the rod. I rotated it under the mocroscope and could celarly see the 'V' shape all the way around. The smooth undanaged groove caught the light more easily and was obvious. The damaged one is scratched around much of the surface, but the undamaged standoff is very smooth. Eric, were both aluminum standoffs in the container with the extra ruby one taken off ETMX, or was one of them new? in any case, see Attachement 4 for a comparison. The believed damage is somewhat visible on the top edge of the lower standoff in the photo. 

[Edit:] Also, in the drawings it looks like the specified radius for the bottom of the groove (0.001 in) is smaller than the radius of the wire (0.00085 in). This would prevent having two clean points of contact like Steve and Gautam were describing as the goal. This is also true of drawings for the new Sapphire guiderods, though the dimensions are in metric units the specified radius of the groove bottom is smaller than the wire's diameter, but larger than its radius. Maybe this providied the initial ability for the wire to move around and carve two distinct grooves. 

Attachment 1: wire_damage_4.jpg
wire_damage_4.jpg
Attachment 2: wire_damage_zoom.jpg
wire_damage_zoom.jpg
Attachment 3: no_wire.jpg
no_wire.jpg
Attachment 4: comparison_2.jpg
comparison_2.jpg
  12347   Thu Jul 28 13:49:44 2016 gautamUpdateGeneralETMX magnets re-reglued
Quote:

[gautam, ericq]

Brief summary, some pictures and such follow in the daytime. 

  • All magnets enclosed by OSEMS were knocked off. (This means one side magnet remained, on the guiderod side)
  • Glue residue cleaned from ETMX
  • Dumbbell feet were cleaned of glue, finely sanded, re-scored with a razor blade
  • The magnet gluing fixture was painstakingly aligned multiple times, adjusting the shim height to get a magnet/wire groove clearance of 0.25-5mm
  • Epoxy mixed, test batch cured perfectly, applied to 5 dumbbells (after triple-checking polarities) and set to glue. 

The epoxy needs at least 12 hours of room temperature air curing, so no touchy until 3:30PM on Jul 28!

Attachment #1 - After multiple trials shimming the magnet gluing rig with teflon spacers, we think that we managed to find a configuration in which the side magnet edge is between 0.25 mm and 0.5 mm from the groove in the ruby wire standoff in which the wire will sit. 

Attachment #2 - Zoomed in view of the side magnet.

Of course we won't know until we suspend the optic, but we believe that we have mitigated the misalignment between the side OSEM axis and side magnet.

The short term plan is to try and suspend ETMY in the end chamber and have a look at the alignment between all magnets and OSEM coils for it. Once the epoxy on ETMX is cured, we will try and suspend the optic again, this time taking extra care while tightening the wire clamps.


Unrelated to this work: Bob just informed me that we had left the air bake oven on overnight - this unfortunately melted the plastic thermocouple inside.

Attachment 1: ETMX_sideMagnet.JPG
ETMX_sideMagnet.JPG
Attachment 2: ETMX_sideMagnet_zoomed.JPG
ETMX_sideMagnet_zoomed.JPG
  12346   Thu Jul 28 08:42:36 2016 SteveUpdateSUS wire clamp

The clamp is in the machine for milling off the grooves. It's condition is normal. The edges needs some 800 sand paper so it is not sharp anywhere.

How did those nicks get on the edge? Fortunatelly they did not aligned with the wire.

 

Attachment 1: e12342.jpg
e12342.jpg
  12345   Thu Jul 28 03:22:28 2016 ericqUpdateGeneralETMX magnets re-reglued

[gautam, ericq]

Brief summary, some pictures and such follow in the daytime. 

  • All magnets enclosed by OSEMS were knocked off. (This means one side magnet remained, on the guiderod side)
  • Glue residue cleaned from ETMX
  • Dumbbell feet were cleaned of glue, finely sanded, re-scored with a razor blade
  • The magnet gluing fixture was painstakingly aligned multiple times, adjusting the shim height to get a magnet/wire groove clearance of 0.25-5mm
  • Epoxy mixed, test batch cured perfectly, applied to 5 dumbbells (after triple-checking polarities) and set to glue. 

The epoxy needs at least 12 hours of room temperature air curing, so no touchy until 3:30PM on Jul 28!

  12344   Wed Jul 27 22:42:00 2016 PrafulUpdateElectronicsEM172 Amplifier

I recreated Den's microphone amplifier circuit on a solderless breadboard to test it and make sure it does what it's supposed to. So far it seems like everything is working- I'll do some testing tomorrow to see what the amplified output is like for some test noises. Here's the circuit diagram that Den made (his elog as well https://nodus.ligo.caltech.edu:8081/40m/6651):

I'm not sure why he set up the circuit the way he did- he has pin 7 grounded and pin 4 going to +12V while in the datasheet for the opamp (http://cds.linear.com/docs/en/datasheet/1677fa.pdf), pin 7 goes to positive voltage and pin 4 goes to negative voltage. There's some other strange things about the circuit that I don't really understand, such as the motivation for using no negative voltage source, but for now I'm going to stick with Den's design and then make some modifications after I have things working and a better understanding of the problem.



Here's my current plan:

-Make sure Den's amplifier works, test it out and make changes if necessary

-Make multiple amplifier circuits on soldering breadboard

-Either make a new amplifier box or reuse Den's old box depending on how many changes I make to the original circuit

-Solder EM172s to BNC connectors, set them up around the floor suspended

-Get the amplifier box hooked up, set up some data channels for the acoustic noise

-Add new acoustic noise tab to the summary pages

 

Den also mentioned that he wanted me to measure the coupling of acoustic noise to DARM.

  12343   Wed Jul 27 16:56:47 2016 ranaMetaphysicsSUSBroken wire

For the rest of this vent, at least, we need to start using the EQ stops more frequently. Whenever the suspension is being worked on clamp the optic. When you need it to be free back off the stops, but only by a few hundred microns - never more than a millimeter.

Best to take our time and use the stops often. With all the magnets being broken off, its not clear now how many partially cracked glue joints we have on dumbells which didn't completely fall off.

  12342   Wed Jul 27 15:03:02 2016 sisyphusMetaphysicsSUSBroken wire

While tightening the bolts on the ETMX wire clamp, the wire broke. All four face magnets broke off. 

Fortunately, no pieces were lost.

  12341   Wed Jul 27 11:40:48 2016 steveSummarySUSoplev laser summary updated

 

Quote:

 

Quote:

 

Quote:

 

Quote:

 

                  2005              ALL oplev servos use Coherent DIODE LASERS # 31-0425-000, 670 nm, 1 mW

    Sep. 28, 2006              optical lever noise budget with DC readout in 40m,  LIGO- T060234-00-R, Reinecke & Rana

    May  22, 2007              BS, SRM & PRM  He Ne 1103P takes over from diode

    May  29, 2007              low RIN He Ne JDSU 1103P selected, 5 purchased sn: T8078254, T8078256, T8078257, T8078258 & T8077178 in Sep. 2007

    Nov  30, 2007               Uniphase 1103P divergence measured

    Nov. 30, 2007               ETMX old Uniphase 1103P  from 2002 dies: .............., running time not known......~3-5 years?

    May 19, 2008               ETMY old Uniphase 1103P from 1999 dies;.....................running time not known.....~    ?

    Oct.  2, 2008                ITMX & ITMY are still diodes, meaning others are converted to 1103P earlier

 

                     JDSU 1103P were replaced as follows:

   May 11, 2011                ETMX replaced, life time 1,258 days  or 3.4 years

   May 13, 2014               ETMX , LT 1,098 days or 3 y

   May 22, 2012               ETMY,  LT 1,464 days or  4 y

   Oct.  5, 2011                BS & PRM, LT 4 years,  laser in place at 1,037 days or 2.8 y

   Sep. 13, 2011               ITMY  old 1103P &    SRM    diode laser replaced by 1125P  ..........old He life time is not known, 1125P in place 1,059 days or 2.9 y

   June 26, 2013              ITMX 622 days or 1.7 y    note: we changed because of beam quality.........................laser in place 420 days or 1.2 y

 

  Sep. 27, 2013               purchased 3 JDSU 1103P lasers, sn: P893516, P893518, P893519 ......2 spares ( also 2 spares of 1125P of 5 mW & larger body )

 

      May  13, 2014             ETMX,  .............laser in place 90 d

      May  22, 2012             ETMY, 

     Oct.  7,  2013             ETMY,  LT  503 d  or  1.4 y............bad beam quality ?

     Aug. 8,  2014              ETMY,  .............laser in place   425 days  or  1.2 y

 

      Sept. 5, 2014              new 1103P, sn P893516  installed at SP table for aLIGO oplev use qualification

     

           May 23, 2016             ITMX dead laser sn P845648 replaced after 1062 days [2.9 yrs] by 1103P, sn P859884, with output output  2.6 mW, nicely round beam quality at 15 meters.

               July 27, 2016             2  new 1103P from Edmonds in: P947034 & P947039, manf. date April 2016,

  12340   Tue Jul 26 18:07:57 2016 gautamUpdateGeneralETMX suspension - Further OSEM Tests

Today, we did the following:

  • Once again, inserted all four face OSEMs till the sensor voltage readouts were approximately half their saturation value. The presence of some ferromagnetic material in the Honeywell components makes this tricky as each coil is coupled to the other three, but we were able to converge to a point where all the voltage readouts were oscillating around a mean value of ~40-60% of their maximum value, with all the damping loops OFF.
  • Turned on all damping loops, and verified that the OSEM positioning was indeed such that the sensor readout is nominally around 50% of the saturation value. The air buffeting around the clean bench means that the damping isn't nearly as effective as it is inside the vacuum chamber.
  • Attempted to increase the gain on the damping loops - we first switched out the Chebyshev low-pass filter in all the damping loops for something a little less aggressive, to allow us to turn up the gain. However, this experiment wasn't a success, when we turned the damping loops on, they were ringing the optic up.
  • At this point, Eric checked the offset sliders (summed in via the slow system) and saw that they were not zero. We zeroed these, but naturally, they destroyed the OSEM positioning equilibrium we had established earlier. So we had to go back and re-position the OSEMs
  • After re-centering the face OSEM magnets relative to the LED-PD pair, we insertd the side OSEM such that the side magnet completely occluded the PD. Interestingly, Eric noticed that the magnetic attraction between OSEM and magnets conspired to center the side magnet fairly well in the side OSEM, when it completely blocked the PD. However, when he returned the side OSEM coil to its nominal operating position of approximately half-blocking the PD, some minor misalignment was re-introduced (i.e. even when the optic was swinging mostly along the axis of the side OSEM, the voltage readout did not quite go down to 0). 
  • We then decided to compare the spectra for the error signals for the 4 DOFs with the current configuration (i.e. suspension clamped down to table top, optic freely hanging, all OSEMs reasonably well centered, and with the ETMX SUS model reverted to its normal state) to some reference (see Attachment #1). I initially thought I would wait for the optic to settle down a bit more before taking the spectra, but it doesn't seem to be showing any signs of getting any quieter in the last one hour. In Attachment #1, I have plotted as reference the spectra of the error signals from the early hours of 4 July 2016, at which point we were at atmosphere but the heavy doors were not yet off, so this is not really a fair comparison, but we don't really have a period in which the optic was exposed to the atmosphere and with the OSEMs in place, at least from this vent. Colors are identical for a given DOF, with todays trace as a solid line, and the reference dashed.
  • We did not check the room available to install some shimming piece of metal in the side OSEM holder, as a possible solution to solve the misalignment problem. Steve has already found pieces of varying thickness, and they are soaking in acetone right now, we plan to air bake them tomorrow. 

I will have another look at the spectra tomorrow morning, to see if the damping improves overnight. 

Attachment 1: OSEM_spectra.pdf
OSEM_spectra.pdf
  12339   Tue Jul 26 17:41:59 2016 KojiSummaryVACPurge compressed air system

We have no number for the CFM without calculation. We can't assume a random number like 10-15

  12338   Tue Jul 26 16:01:32 2016 SteveSummaryVACPurge compressed air system

Thanks for checking this out Koji

The builder in 1996 was Process System International, Inc ( Westborough, MA ) It does not exist any longer or I just could not find them. Flow diagramm at Atm1

Should I be keep looking for a company who could quote us for building a similar smaller unit with 10 - 15 cfm flowrate?

Note: my intension with the two mobile-overhead HEPA filter was the same as John Worden's " clean air overpressured tent " at chamber entrance.

Atm2, Our unit has 650 cfm, velocity 90 fpm at resistance 0.5"    It may be enough to give a little overpressure if we seal it well to the chamber

We use to use them to minimize dirt getting inside the chanbers.

Attachment 1: cleanAirSupl.pdf
cleanAirSupl.pdf
Attachment 2: mobileHEPA.jpg
mobileHEPA.jpg
  12337   Tue Jul 26 14:24:38 2016 KojiSummaryVACPurge compressed air system at LHO

I've visited the purge clean air system at LHO Yarm mid-station with John Worden.

The system is described C981637. There is a schematic in C981637-06-V (Vol.6).pdf although the schematic has some differences (or uncorrected mistakes).

This system is intended to provide positive pressure when a soft cover is attached to a chamber door. When the door is open, the purging does not help to keep the chamber clean because the flow is too slow. This protection has to be done with overhead HEPA filters (22x5000cfm). It may be possible that this purge air helps the tube not to allow dusts to come in. But before using this, the chambers and the tubes have to be cleaned, according to John.


- Here at the site, the purge air system is started up a day before the vent. This system is used for the vent air, the purge air, and turbo foreline filling.

- Air intake (attachment 1): At the site, the air is intaken from the VEA. We want to incorporate somewhat clean air instead of dirty, dusty, outside air.

- Initial filter (attachment 2): a high volume filter before the compressors.

- The compressors (attachment 3, 4) are 5x 6 horse power air compressor each goes up to 160 psi. They are turned on and off depending on the demand of the air. Which is turned on is revolved by the controller to equalize the compressor usage hours.

- The compressed air goes through the air cooler (heat exchanger) to remove the heat by the compressor work.

- This air goes through prefilters and accumulated in the air receiver (100psi) (attachment 5). This receiver tank has an automated vent valve for periodical water drainage at the bottom.

- The accumulated air is discharged to twin drier towers (attachment 6, blue). The tower is operated by the controller (attachment 7) alternately with a period of 4min (or 10min by setting). When one of the towers is working, a humid air comes from the bottom and the dry air is discharged from the top. A part of the dry air goes into the other tower from the top to the bottom and dries the tower. There is a vent at the bottom to discharge water periodically.

- The dried air goes through 4 types of filters. After the last filter, all of the plumbing should be made of stainless steel to keep cleanliness.

- The air goes to the pressure reducing regulator (attachment 8, gray). The final flow speed at the chamber side is 50cfm max, according to John.

- The lower pressure air goes through the final filter (attachment 8, blue). As the pressure is low, this filter is big in order to keep the volume of the air flow.

- The purge air is supplied to the chamber side with KF50 (attachment 9). There is a vent valve (attachment 10) for safety and also to run a dry air for at least a day before the use to clean up the supply line. The purge line is disconnected when no in use.

- The entire system (attachment 11) and size comparison (attachment 12).

 

Attachment 1: air_intake.jpg
air_intake.jpg
Attachment 2: initial_filter.jpg
initial_filter.jpg
Attachment 3: compressors1.jpg
compressors1.jpg
Attachment 4: compressors2.jpg
compressors2.jpg
Attachment 5: air_receiver_dryer.jpg
air_receiver_dryer.jpg
Attachment 6: drier.jpg
drier.jpg
Attachment 7: drier_controller2.jpg
drier_controller2.jpg
Attachment 8: pressure_regulator_and_last_filter.jpg
pressure_regulator_and_last_filter.jpg
Attachment 9: chamber_side_supply.jpg
chamber_side_supply.jpg
Attachment 10: vent_valve_for_line_cleaning.jpg
vent_valve_for_line_cleaning.jpg
Attachment 11: the_whole_system.jpg
the_whole_system.jpg
Attachment 12: size_comparison.jpg
size_comparison.jpg
  12336   Tue Jul 26 09:56:34 2016 ericqUpdateCDSc1susaux restarted

c1susaux (which controls watchdogs and alignments for all non-ETM optics) was down, the last BURT was done yesterday around 2PM. 

I restarted via keying the crate. I restored the BURT snapshot from yesterday.

  12335   Mon Jul 25 20:37:46 2016 gautamUpdateGeneralETMX suspension - Side OSEM test

[Lydia, gautam]

Summary: We did some preliminary tests to check if at least one of the side magnet positions is usable for the side OSEM. We mainly wanted to check how much dynamic range we lose because of the sub-optimal longitudinal positioning of the side magnet. We found that when the side magnet was mainly moving along the axis of the side OSEM (with minimal yaw motion as gauged by eye), the PD voltage bottomed out at ~80 counts (while the completely unoccluded readout was ~800 counts). 

Details:

  1. First, we placed the face OSEMs into their holders one by one, and adjusted their position till the readout was approximately half the saturation value (as judged by the average value of the readout, at this point, the mirror was still swinging around a fair bit).
  2. Next, we enabled the POS, PIT and YAW damping (with all existing settings unchanged), but with the SD coil input and output disabled. We had to increase the watchdog threshold to ~600mV.
  3. Once the optic was reasonably well damped (~70mV on the watchdog was the best we saw), I put in the side OSEM till the PD was completely occluded. At this point, I enganed the earthquake stops, and then released the mirror such that it was freely hanging. I then observed the optic by eye, and noted a time when the dominant motion was along the axis of the side OSEM coil (i.e. minimal YAW motion). 
  4. Attachment #1 shows time series plots of the 5 OSEM PD voltage monitors. Perhaps because the side OSEM input was disabled, the damping wasn't as efficient as it normally is (also there is a fan blowing air around the clean bench). But at the point indicated on the plot, the YAW motion was negligible to the eye, while the dominant motion was along the axis of the coil. During this time, the readout bottomed out at approximately 10% of the saturation value (towards the end of these plots, I disabled the damping loops and began pulling the OSEMs out one by one). Because the damping was imperfect, this is only an approximate guess of how much dynamic range we are losing. But does this warrant regluing the side magnet?

Other remarks:

  • The 4 face magents were reasonably well centered in the coil. While Eric and I were looking at this earlier today, the LL magnet looked a little close to the coil, but after putting all 4 OSEMs in, the situation looked reasonable to the eye. I couldn't take pictures because of space constraints, and furthermore, it's almost impossible to hold the camera in the correct vertical position. 
  • Steve, Eric and I couldn't find the OSEM gender changer anywhere in the lab and it wasn't in the box it was advertised to be in. So we made a custom cheater cable, and cleaned it by wiping with Isoprop., and wrapped it in foil for use in this test. The OSEM pins should probably be cleaned before we put these back in vacuum. 
Attachment 1: osem_test_2016_07_25.png
osem_test_2016_07_25.png
  12334   Mon Jul 25 17:01:54 2016 SteveUpdateSUSETMX suspension - progress

ITMY side  : Magnet od 1.9 mm so wire to magnet gap ~ 0.2-0.3 mm

 

Attachment 1: ITMYside2011.JPG
ITMYside2011.JPG
  12332   Mon Jul 25 13:47:19 2016 ranaUpdateGeneralETMX suspension - progress

Ugh. frown

Steve, please look into getting some plated magnets (either SmCo or NdFeB is OK) of this size so that we can install cleaner magnets by the next vent.

  12331   Mon Jul 25 13:01:49 2016 ericqUpdateGeneralETMX suspension - progress

(Full resolution versions of the photos in this ELOG are on picasa)

The OSEM gender changers were not in the box labelled as such, we need these to be able to use the OSEMS to see just how bad the side magnet alignment is, and to do any kind of damping for the fine pitch balancing. The hunt is on.

In the meantime, Gautam and I checked out the standoff seating, and alignment of the face OSEMS (after slightly adjusting the wire length - I guess some sagging is still happening).

With a bit of poking, we convinced ourselves that we sat the standoff in contact with the optic's barrel. Amazingly, we were able to maintain the coarse pitch balance of the optic.

We then partially inserted the face OSEMS, to check their magnet alignment. ("partially" means that the OSEM is not actually enclosing the magnet, we don't want to knock anything off) They seem ok, but not perfect. These magnets were not removed or reglued, so presumably their alignments should be unchanged.

  12330   Mon Jul 25 12:22:18 2016 ranaUpdateComputer Scripts / ProgramsFinished MEDM Tab on Summary Pages

Looks pretty great. However, there's two problems:

1) Some of the MEDM screens don't show the time. You can fix this by editing the screens and copy/paste from screens which have working screens.

2) The snapshot script seems to not grab the full MEDM screen sometimes.

These are not a very big deal, so you can get the microphones working first and we can take care of this afterwards.

  12329   Mon Jul 25 10:54:55 2016 PrafulUpdateComputer Scripts / ProgramsFinished MEDM Tab on Summary Pages

The MEDM screen capture tab is now working and up on the summary pages: https://nodus.ligo.caltech.edu:30889/detcharsummary/day/20160725/medm/

Please let me know if you have any suggestions or notice any issues.

  12328   Sun Jul 24 03:43:56 2016 ericqUpdateGeneralMagnet positioning

When Koji and I were gluing magnets to ETMY, we decided to position the side magnet based on the empirically observed offsets from the standoff groove seen at other side magnet locations. Specifically, we figured that the magnet should be glued 1.25mm closer to the HR surface than the wire groove.

However, Steve has told me that he believed that this distance should be something like 0.5mm. 

I used the 1.25mm figure when gluing the ETMX side magnets, which now do not align well to their OSEM mounts. While it is certainly possible that I made an error when shimming the fixture, I think it is also possible that this figure was incorrect. 

Sadly, after poring through the DCC and various elogs, I have not been able to come up with a definitive answer on what this offset should actually be.

One approach is to examine the suspension tower dimensions. I.e. when subject only to gravity, the wire loop should lie in the plane of the back face of the top block of the suspension, as it is constrained by the clamps. Thus, the standoff grooves also lie in this plane. The center of the side OSEM mounting holes are about 1.64mm in front of this plane, which is larger than the 1.25mm figure that Koji and I came up with. Examining the picture Gautam posted of the marginal magnet/OSEM alignement, we see that this figure would in fact move the magnet in the wrong direction...

ELOGs in which the intitial side magnet gluing and fixture shimming are detailed do not reference the absolute position of the side magnet, nor do they include any pictures of their fixture setup. (Some links for the curious: 2652 2654 2668)

The DCC isn't much help either, as it is not clear what version of the gluing fixture we actually have. There is a drawing for a 40m specific version, but it includes swappable side-magnet-pickle-picker-slots to achieve different positions for different (circa 2001) optics; this is not the kind of fixture we currently have in our possesion. (https://dcc.ligo.org/D010131) I have discovered that some versions of this fixture (https://dcc.ligo.org/d990168) include an assumed 0.5deg wedge angle and thus position the two side slots differently. Although the fixture we have has no identifying marks on it whatsoever (naturally), I measured the two side slots to be different in axial position by roughly 0.6mm, which is consistent with a 0.5deg wedge. Furthermore, the sign of this difference indicates that this fixture ring is designed for the opposite wedge orientation than our ETMs, which have a 2.5deg wedge, making this fixture wrong by 3deg (which is ~4mm over the diameter of the optic). 

We did not account for this for either ETMX or ETMY, so this is another source of error, but this does not give us much guidance on what the real absolute magnet position should be.

  12327   Fri Jul 22 23:03:23 2016 gautamUpdateGeneralETMX suspension - progress

[ericq, gautam]

Summary:

Today, we attempted to progress as far as we could towards getting the mirror suspended and gluing the second wire standoff. We think we have a workable setup now. At this stage, the suspension wire has been looped around the magnet, the second wire standoff has been inserted, coarse pitch balancing has been done, and we have verified that side OSEM/magnet positioning is tenable. Details below.

Details:

  1. First we verified that the epoxy on the side magnet re-glued yesterday had dried (verified using control setup of epoxy in aluminum foil + copper wire - we didn't perform any further tests like pulling the magnet off the tabletop as we were satisfied)
  2. We placed the optic inside the suspension cage, resting on the 4 lower earthquake stops.
  3. We looped the suspension wire around the optic. This is a somewhat challenging procedure. After consulting the documentation, we decided to follow the given advice and loop the wire around from the bottom of the optic, one side at a time. It is tricky to thread the wire between the two lower earthquake stops and get it up around the side. The side magnets were an unexpected ally in this effort as they served as some sort of intermediate checkpoint from which we could pull the wire further up. We then lightly clamped it to the winches mounted atop the suspension cage.
  4. After verifying that we had routed the wire correctly through the various stages (primary and secondary suspension points at the top of the cage), we placed the wire under very slight tension by had, and then tightly clamped the wires in the winches (we then cut off the excess length).
  5. In this state, we proceeded to install the second wire standoff (having verified that the wire was indeed sitting in the groove on the other side). 
  6. We then proceeded to raise the optic to the desired height (center of optic to 5.5 inches above the table top) with the help of the microscope and the lines on the barrel side. 
  7. Next, we attempted to freely suspend the optic (i.e. no contact with the viton tips). We were initially unsuccessful but Eric did some fine adjustment of the (unglued) standoff to achieve a stable configuration. However, the wire is now really close to the magnet - although it is not clear to me if it is touching the magnet as we initially suspected - see Attachment #1, it may be that if the wire is touching something, it is the dumbbell and not the magnet itself. While this is clearly not ideal, we think that this setup is workable as is. If after doing the pitch balancing, if the deviation of the wire becomes much more pronounced, we may have to re-glue the side magnet. In any case, both the horizontal scribed lines are now 5.5 inches above the table top.
  8. We then brought over the OSEMs from the ETMX vacuum chamber to the cleanroom. As a first check, we wanted to ensure that one of the side magnets could accommodate an OSEM (because both side magnets have been re-attached after the optic was removed from the old suspension). Attachment #2 suggests that this is possible, even though the relative positions of the side magnet and the shadow sensor may be sub-optimal. We will only really know after hooking up the electronics.

Remarks:

  1. We found that after a few hours, there was some sag introduced in the wire, presumably it stretched into an equilibrium position under the weight of the optic. We will re-check the heights tomorrow while conducting further tests.

Immediate to-do:

  1. Insert all OSEMS. Ensure that the magnet positions relative to the coil are compatible.
  2. Enable damping loops. We have a cable coming from the IFO area into the cleanroom through a hole-in-the-wall. We are missing a DB25 gender changer at the moment. 
  3. Do the pitch balancing.
  4. Glue the second standoff in place. 

Other attachments:

Attachment #3 - Unglued stand off with wire in the groove, mirror freely suspended.

Attachment #4 - Glued stand off with wire in the groove, mirror freely suspended. Clearance between wire and magnet looks reasonable.

Attachment #5 - Barrel of optic (underside), mirror freely suspended. The wire seems to be in a reasonable orientation along the barrel, albeit not perfectly parallel.


Koji just pointed out that we should check that the unglued ruby standoff is in good contact with the barrel of the optic. Attachment #1 suggests that maybe this is not the case. If you zoom into Attachment #1, it is not clear if the standoff is sitting on the glue.

Attachment 1: IMG_2828.JPG
IMG_2828.JPG
Attachment 2: IMG_2826.JPG
IMG_2826.JPG
Attachment 3: IMG_2821.JPG
IMG_2821.JPG
Attachment 4: IMG_2822.JPG
IMG_2822.JPG
Attachment 5: IMG_2823.JPG
IMG_2823.JPG
  12326   Fri Jul 22 05:20:26 2016 AakashUpdateGeneralTemperature measurements across the enclosure | SURF 2016

Please find the new attached plots and the new script.

Attachment 1: coherence.pdf
coherence.pdf
Attachment 2: transferfunc.pdf
transferfunc.pdf
Attachment 3: transferfuncdB.pdf
transferfuncdB.pdf
Attachment 4: script22jul2016.m
filename='2315on5july.dat';
data=importdata(filename);
%temperature data outside the enclosure on channel 2
data1=data(:,2); 
%temperature data inside the enclosure on channel 3
data2=data(:,3); 

%sampling frequency in Hz
fs=100; 
... 40 more lines ...
  12325   Fri Jul 22 03:02:37 2016 KojiUpdateCOCFC painting

[Koji Gautam]

We have worked on the FC painting on ITMX and ITMY. We also replaced the OSEM fixing screws with the ones with a hex knob.
This was done except for the SD OSEM as the new screw was not long enough. We left an allen-key version of the screw for the SD OSEM.

All the full-resolution photos can be found on g-photo.


ITMY

Attachment1: The barrel was pretty dusty. Some dusts were observed on the HR face but it was not so terrible. The barrel and the HR face were blown with the ionized N2 and then wiped with IPA. The face wiping was done n a similar way as the drag wiping.

Attachment2: FC was applied to the HR surface.

Attachment3: The AR surface was also painted with FC. The brush touched the coil holder.

Attachment4: The brush touched the coil holder. Another PEEK tab was applied to remove this FC stain on the metal holder.

Attachment5: This is the result of successful removal of the FC stain.

ITMX

Attachment6: The OSEM arrangement before removal. We confirmed that the OSEM arrangement was as described on Wiki.

Attachment7/8: The ITMX was obviously a lot dirtier than ITMY. The barrel accumulated dusts.

Attachment9: This is the HR face picture with large dusts on it.

Attachment10: The HR surface was painted with FC.

Attachment11: This is the AR surface with FC painted.

Attachment 1: ITMY_barrel_dust.jpg
ITMY_barrel_dust.jpg
Attachment 2: ITMY_HR_FC.jpg
ITMY_HR_FC.jpg
Attachment 3: ITMY_AR_FC.jpg
ITMY_AR_FC.jpg
Attachment 4: ITMY_drip_removal.jpg
ITMY_drip_removal.jpg
Attachment 5: ITMY_drip_removed.jpg
ITMY_drip_removed.jpg
Attachment 6: ITMX_OSEMS.jpg
ITMX_OSEMS.jpg
Attachment 7: ITMX_barrel_dust1.jpg
ITMX_barrel_dust1.jpg
Attachment 8: ITMX_barrel_dust2.jpg
ITMX_barrel_dust2.jpg
Attachment 9: ITMX_HR_dusty.jpg
ITMX_HR_dusty.jpg
Attachment 10: ITMX_HR_FC.jpg
ITMX_HR_FC.jpg
Attachment 11: ITMX_AR_FC.jpg
ITMX_AR_FC.jpg
  12324   Thu Jul 21 22:02:35 2016 varunUpdateCDSDAFI update: Frequency warping

The code for frequency warping contained a "printf()" command, which had caused the system to crash in one another instance (refer elog 12320) . Hence, I tried running the code tody by removing this line. Unfortunately, this did not work. the model still crashed. Attached is the screenshot of the FE status.

Attachment 1: 07212016.png
07212016.png
  12323   Thu Jul 21 21:38:44 2016 AakashUpdateGeneralTemperature measurements across the enclosure | SURF 2016

I have made the changes as suggested by Gautam.

  12322   Thu Jul 21 17:01:59 2016 ericqUpdateGeneralETMX magnet gluing/guiderod excess glue removed

One step forward, two steps back...

While attempting to suspend ETMX, I broke off a side magnet crying

It is now gluing

(This is *not* the one that was previously glued. I.e., now both ETMX side magnets have been reglued)

  12321   Thu Jul 21 15:03:13 2016 varunUpdateCDSDAFI Update

1) I have added the status summary of the DAFI block to the main FE status overview screen in the c1lsc cloumn. (attachment 1)

2) I have edited all the kissel matrix buttons appropriately, and given them appropriate lables. (attachment 2)

Attachment 1: festatus.png
festatus.png
Attachment 2: matrices.png
matrices.png
  12320   Thu Jul 21 14:27:24 2016 varunUpdateCDSDAFI Update: Arbitrary Math block

Summary: I have added an arbitrary math block to the DAFI model, which takes two inputs, say X and Y, and can perform various unary and binary operations on them:

Details:

  • Unary Operations:
  1. Delay - There exists a text-based input to specify the amount of delay to be given to a particular signal.
  2. sin()
  3. cos()
  • Binary Operations:
  1. Weighted addition and multiplication: The output is calculated according to the relation: A*X + B*Y + C*X*Y. A, B, C are constant inputs, which can be given through text-based inputs in the GUI.
  2. MAX{X,Y}
  3. MIN{X,Y}

Attachment 1 shows the existing DAFI gui, updated with cascading of various DSP blocks, upto three levels, button-based ENABLE and DISABLE controls for all blocks except arb. math (the control on arb. math. is achieved by clicking on the block.) On clicking the arb. math block one is taken to the dedicated arb. math screen, which has enable buttons for all the processes listed above. A screenshot of this screen is in attachment 2. There is one control input, which controls all the unary operations on X and the binary operations on X and Y, and another control input which controls the unary operations on Y. switching on a particular arb. math process gives a particular control input, which choses the appropriate section of the code. At a time, only one process from the top grey block (corresponding to unary operations on X and binary operations on X and Y) and one process from the bottom grey block (corresponding to unary operations on Y) can be selected. Thus, the outputs which go from the arb. math block to the intermediate matrices (MATRIX1L or MATRIX2L) are:

a) Either an output of unary operation on X or a binary operation on X and Y, the specific one depending upon the control input,

    and

b) Output of a unary operation on Y, again the specific one depending upon the control input

Thus there is apparent asymmetricity in the action of the arb. math block on the two inputs. However, this is done in order to reduce to total number of outputs and control signals, and this can be easily taken care of by interchanging the inputs before the block. 

While compiling this code, the c1lsc machine had crashed once, it was found that this was due to a stray "printf()" command in the c code. This glich in the code now stands rectified  There is a possibility that the previous incidents of the code crashing could also be due to the existence of a printf() command. 

Preliminary Testing: I have done a preliminary testing of the arb math block, i.e. verified that on enabling the sin and cos processes, the output is less that 1, on swithching on the process of weighted avarage and multiplication, the output looks like it is right, for a few simple values of A, B, C, like 0, 1, etc. The delay block however is giving zero output for delay of more than 6 samples.

 

 

 

 

Attachment 1: dafioverview.png
dafioverview.png
Attachment 2: arbmath.png
arbmath.png
  12319   Thu Jul 21 12:03:35 2016 varunUpdateCDSDAFI update: Humming noise in DAFI output

 

Summary: There was always a constant humming noise in the output of speakers of both the audio channels. Tried to resolve the problem. Details are given below:

Details: The source of the noise was the typical 60 Hz (and harmonics), ~13 mV peak to peak output, in at least three channels of the DAC. (two coming from the DAF module, and one not related to the DAF.) Attachment 1 shows the noise in both the DAF channels. As compared to that, the signal coming through the AGC weak, about 6 mV RMS, about the same order as noise. In order to resolve this, the gain of the AGC was increased, so that the RMS output voltage of the Fibox (FBAO, the one at the output) was about 1.23 V RMS. It is approximately equal to +4 dBu, which is the typical expected output of the Fibox, according to the datasheet.

Attachment 1: New_Doc_13.pdf
New_Doc_13.pdf New_Doc_13.pdf
  12318   Thu Jul 21 08:24:20 2016 SteveUpdatePEMants alart

The heat drives the ants to the lab!  Make sure the light doors are tight on the chambers.

 

Attachment 1: ants_Alart.jpg
ants_Alart.jpg
  12317   Thu Jul 21 05:22:26 2016 AakashUpdateGeneralTemperature measurements across the enclosure | SURF 2016

I have measured the transfer function of temperature fluctuations inside the enclosure to that of the temperature fluctuations outside. The transfer function has been estimated by using 'tfestimate' which is library function in Matlab and which estimates the transfer function based on Welch's method. The attached plots shows the transfer function of the temperature inside the enclosure to that of outside temperature.

fig1.pdf

fig2.pdf

In order to determine a relation between temperature inside the enclosure to that of the outside temperature, I have calculated the mean squared coherence.  I have used Matlab's 'mscohere' library function which uses Welch's method to calculate the coherence. Attached plot shows the coherence between the temperature across the enclosure.

fig3.pdf

Also, I have attached the matlab script which I used for generating these plots.

script21jul2016.m

Attachment 1: script21jul2016.m
filename='2315on5july.dat';
data=importdata(filename);
%temperature data outside the enclosure on channel 2
data1=data(:,2); 
%temperature data inside the enclosure on channel 3
data2=data(:,3); 

%sampling frequency in Hz
fs=100; 
... 30 more lines ...
Attachment 2: fig1.pdf
fig1.pdf
Attachment 3: fig2.pdf
fig2.pdf
Attachment 4: fig3.pdf
fig3.pdf
  12316   Wed Jul 20 14:03:01 2016 ericqUpdateSUSEffect of thicker ruby standoffs

The new ETMX ruby guide rods are slightly thicker than the old aluminum ones; specifically 1.27mm vs 1.0mm.

Since we did not change the guide rod location in response to this fact, the vertical position of the suspension point changes, which in turn changes the dynamics of the suspension. Specifically, since the standoff is placed below the guide rod, the suspension point is lowered, which makes the pitch mode softer. I crunched a few numbers and have determined that this effect should not be a problem.

Given the wiki's value of the ETMX pitch resonance frequency of 0.829 Hz, I predict a the new pitch resonance frequency of 0.800 Hz.

(wiki link: https://wiki-40m.ligo.caltech.edu/Suspensions/Mechanical_Resonances)


A useful document about the dynamics of our suspension can be found at T000134

From this document, one will find that the effect of changing the suspension point height over the optic center of mass,`b`, on the pitch resonance frequency (while keeping all other dimensions equal) to be:

\frac{\omega'_\theta}{\omega_\theta} \approx \sqrt{\frac{b'}{b}}

The top of the standoff is fixed by the guide rod, so let's say that b' is given by the position of the center of the Ruby standoff. This is then smaller than the previous b by the differences in the radii of the standoffs:

b' = b - \frac{\oslash_{\textrm{ruby}}-\oslash_{\textrm{Al}}}{2}

The nominal value of b is 0.985mm. Thus, the pitch resonance frequency is changed by factor of 0.965, i.e. 3.5% smaller. Then, taking the wiki value of 0.829 Hz results in 0.800Hz, a 30mHz decrease.

  12315   Wed Jul 20 13:58:55 2016 SteveUpdateSUSSUS damping out of vac chamber

Cheater cable to be used in clean room pitch gluing alingment.

Satelite amp needs to be there.

Atm 2-3, The ETMs  suspension damping cable  are connected at the end racks. All others go to 1X5

Atm 4-5, The other end of this cable in the high cable tray at 1X3 as shown. We'll disconnect the shorty and move the end to ETMX ( or any sus at 1X5 )

 

Attachment 1: ETMXrack.jpg
ETMXrack.jpg
Attachment 2: fromETMX-satBox.jpg
fromETMX-satBox.jpg
Attachment 3: susDampingSatboxCab.jpg
susDampingSatboxCab.jpg
Attachment 4: susDampingSatboxCabl.jpg
susDampingSatboxCabl.jpg
  12314   Wed Jul 20 11:39:07 2016 SteveUpdateVACvent 78 day 20

RGA background scan

Quote:

Vacuum Status: Chamber Open

All chamber annuloses are vented.  Vac Monitor screen is not communicating with gauges. The valve position indicator are working.

RGA is pumped by Maglev through VM2

 

 

Attachment 1: vent78-d20.png
vent78-d20.png
  12313   Tue Jul 19 16:39:29 2016 LydiaUpdateGeneralETMX magnet gluing/guiderod excess glue removed

[ericq, Lydia]

The epoxy arrived. Eric managed to remove the excess glue below the guiderod with a razor blade (see attachment 1). The magnet and dumbell that came apart were reglued successfully and passed the stregth test of picking up the magnet from the table by the dumbell, so the magnet was glued back on the optic and is setting in the gluing apparatus (see attachment 2).

We double checked the polarity against the side magnets on ETMY. Because of the gluing position strategy (a fixed distance toward the HR side from the groove location), the other side magnet appears slightly below the center of the gluing barrel, which after some discussion with Koji was determined to be ok. 

Attachment 1: P7190201.JPG
P7190201.JPG
Attachment 2: P7190203.JPG
P7190203.JPG
  12312   Tue Jul 19 15:30:08 2016 SteveUpdateGeneralHEPA dryer and dust

I set up a simple HEPA filter dryer to dry your clean room garment before  you can put it away into your storage box.

Our lab is dusty ! This is  specially important when we are vented. Please wipe things daily and cover item with foils .... etc.

Attachment 1: HEPAdryer.jpg
HEPAdryer.jpg
  12311   Tue Jul 19 15:16:43 2016 SteveUpdatesafetysafety training

Our new graduate student Lydia received 40m specific safety training.

  12310   Tue Jul 19 13:21:42 2016 JohannesUpdateGeneralVent progress - ETMY attempted repositioning

[Lydia, Johannes]

We moved ITMY from its original position to a place near the access point. We took the OSEMs off first, and noticed that the short flat head screw driver was still a little too long to properly reach the set screws for the lower OSEMs. We were able to gradually loosen them, though and thus remove the lower OSEMs as well. We had to move a cable tower out of the way, but used clamps to mark its position. After making sure the optic is held by its earthquake stops, we moved it to its cleaning location. All magnets are still attached.

  12309   Mon Jul 18 18:44:52 2016 varunUpdateCDSc1lsc FE recovered

c1lsc FE is up and running.

Details:

2) The machine was manually rebooted.

3) c1daf was recompiled and installed, with the problematic piece of code removed.

4) NTP timing was adjusted.

5) Frame Builder was restarted.

6) All models on c1lsc machine were restarted.

Attachment 1 shows the CDS status after the recovery. I wont be trying to run frequency warping immediately, I will first finish implementing the other harmless modules first.

Attachment 1: CDS_status160718.png
CDS_status160718.png
  12308   Mon Jul 18 05:05:05 2016 AakashUpdateGeneralHeater for Seismometer Enclosure | SURF 2016

I took off the silicon rubber heaters which were used by a SURF last year for heating the enclosure. The heater data sheet has mentioned the power dentsities, but I doubted the values. So I wanted to measure the actual power density by these heaters. I think the rubber heaters are broken somewhere within, the surface is not heated evenly. Although I don't have a good quantative reason to use, I was thinking to use a thermoelectric cooling module for the enclosure. 

From the data I collected few days back, I am trying to obtain a transfer function of temperature inside the enclosure to that of outside. My aim is to measure the pole frequency of temperature fluctuations inside the enclosure relative to the outside fluctuations.

 

Attachment 1: HEATERS.jpeg
HEATERS.jpeg
  12307   Sat Jul 16 00:30:42 2016 varunUpdateCDSDAFI update: Frequency warping | c1lsc unresponsive

Summary: I am trying to implement frequency warping/pitch shifting on the real time FE. Here is a description long overdue:

Description: The overall idea is as follows:

The DFT of a frame x_i[n]  is given by X_i[k] = \sum_{n=0}^{m-1}x_i[n]e^{-j2\pi \frac{kn}{m}}. A matrix W containing all W_{kn} = e^{-j2\pi \frac{kn}{m}} for k, n = 1, 2, ..., m can be calculated and predefined in the code. The input arrival rate is 16384 Hz, i.e. once in every 60 $\mu$s time window. Hence, the fourier coefficients can be updated cumulatively in each cycle using the current value of the input, previous value of the fourier coefficient and the components of the W matrix. This will distribute the computational load of the FFT into all the time windows. Similar operations can be carried out for the inverse STFT.

I have written and run a pseudo-real time code on my CPU. The following is the essence:
 Let the frame-length be M, and the intended scale factor of the frequency warping be 'r'. The frame overlap is 50%. At each clock cycle, the following tasks are performed:  (1 to 3 are routine tasks performed at every clock cycle, 4 is a special task performed only when a frame is filled.)
 1) Take input and apply hanning window to it.
 2) Cumulate X_i[k] for every k using the value of x_i[n] (the input) at that particular instant. Also start to cumulate X_{i+1}[k], which will be later transfered to X_i[k].
 3) Because of 4), we now have 'r+1' filled frames corresponding to output fft. Now take the ifft using two consecutive frames corresponding to only two time series points. The computations required for this task are the same as the computations required for calculation of the fourier coefficients iteratively, since the entire time series ifft is not computed.
 4) Do these special tasks after each frame gets filled:
      At this point, the ffts of the current frame and a previous frame is ready. Let us call them X1 and X2.
      Calculate phase difference between the two.
      Calculate all the interpolated |Y_i| in between these two frames depending upon the scale factor.
      Assign phase of X1 to first Y frame and assign increasing phase to all the other Y frames.
      and also do all the usual non-special tasks.

This code takes about 9-10 microseconds for a cycle with special tasks, and 5-6 microseconds for a cycle with routine tasks on my laptop (brought down from 100 microseconds peak time in the earlier offline implementation due to elemination of explicit dft and reduction in fft size), for a frame size of 32 samples. However, when fed into the c1lsc FE, it crashes, as it has done once again today evening, in the same fashon as yesterday. There could be 2 possible reasons:

1) Size of the array containing the W_{kn} matrix elements is too large for the FE memory,

2) the computations are taking up more than 60 microseconds.

Since there are already a few codes with similar array sizes, I am more inclined to think that 2) is more likely.

Another problem that I am anticipating is that for a 32 point dft and a sampling rate of 16kHz, the frequency resolution achieved is about 500 Hz, which is not sufficient if we need to represent seismic signals. The only way I can think of, for representing such signals with a small number of fft points, is to reduce the effective sampling rate, i.e. do DSP on inputs at a much lower rate than 16kHz (say 1kHz, which will give a resolution of ~30 Hz, or 2kHz giving a resolution of ~60Hz). Another advantage of this method is that it frees up more clock cycles for computation, thus the computational load can be further distributed.  The problem in this implementation is that it will increase the delays.

  12306   Fri Jul 15 17:44:37 2016 AakashSummaryGeneralAcromag Setup | SURF2016

Aidan has described the physical connections and initial setup here :  https://nodus.ligo.caltech.edu:30889/ATFWiki/doku.php?id=main:resources:computing:acromag#recovering_from_a_terminal_power_communication_outage  .

Since I used a Raspberry Pi(domenica.martian) for communicating to Acromag(acroey.martian) card, I had to recompile everything for linux-arm architecture. 

For EPICS installation, download the EPICS base from http://www.aps.anl.gov/epics/download/base/baseR3.14.12.3.tar.gz . Installing dependencies, build, install epics at /usr/local/epics. By downloading modbusApp source from https://llocds.ligo-la.caltech.edu/daq/software/source/epics-3.14.12.2_long-source.tar.gz  , build the modbusApp for linux-arm architecture in modules/modbus directory inside epics base.

Put all the files mentioned by Aidan and run a tmux session to grab channels.

Also, pyModbus can be used to read the channels. I'll put the physical connections schematic shortly.

  12305   Fri Jul 15 16:23:51 2016 ericq UpdateGeneralGluing setbacks

[ericq, Lydia]

Here is a picture of the ETMX guide rod post-gluing. There is unfortunately a fair amount of excess. The "tab" is the result from the epoxy travelling along the finger of the fixture arm that held the guide rod.

We set out to glue the previously remove ETMX side magnet, and set up the fixture to do so. For ETMX we needed 3 mm of shimming on the thick side, and 6mm on the thin side.

However, while cleaning the magnet+dumbbell base of epoxy residue, I broke the dumbbell off of the magnet no

We then fetched the spare side magnet that Steve had been holding onto. While cleaning it, it was dropped and dissapeared from this plane of existence nono

So, instead of gluing a side magnet today, we are gluing the existing magnet and dumbbell back together:

Sadly, this used up the last of our EP30.nonono

Though Koji had the foresight to order more(yes), it will not arrive until Monday/Tuesday, so no side magnet gluing until then.frown

  12304   Fri Jul 15 12:21:28 2016 varunUpdateCDSc1lsc FE unresponsive

c1lsc is up and running, Eric did a manual reboot today.

Quote:

Today, at around 10:30, c1lsc machine froze and stopped responding to ping and ssh after I compiled and restarted c1daf. I think it is due to a large array in one of my codes. The daqd.log file shows the following:


..................................................................
CA.Client.Exception...............................................
    Warning: "Virtual circuit unresponsive"
    Context: "c1lsc.martian.113.168.192.in-addr.arpa:5064"
    Source File: ../tcpiiu.cpp line 945
    Current Time: Thu Jul 14 2016 22:27:42.102649102
..................................................................

I think the c1lsc FE may need a hard reboot.

 

  12303   Thu Jul 14 23:38:59 2016 varunUpdateCDSc1lsc FE unresponsive

Today, at around 10:30, c1lsc machine froze and stopped responding to ping and ssh after I compiled and restarted c1daf. I think it is due to a large array in one of my codes. The daqd.log file shows the following:


..................................................................
CA.Client.Exception...............................................
    Warning: "Virtual circuit unresponsive"
    Context: "c1lsc.martian.113.168.192.in-addr.arpa:5064"
    Source File: ../tcpiiu.cpp line 945
    Current Time: Thu Jul 14 2016 22:27:42.102649102
..................................................................

I think the c1lsc FE may need a hard reboot.

  12302   Thu Jul 14 22:29:57 2016 ericq UpdateGeneralETMX guide rod gluing / ETMY Magnet gluing

The pickle puckers came off ETMY cleanly yes ETMY now rests in the ring holder, under a glass jar, with all of its magnets.

We removed the guiderod gluing fixture from ETMX without any apparent damage to the fixture arm, optic, or guiderod epoxy joint. 

I started measuring some distances on the optic for the side magnet gluing, but am not sure of it yet. So, I didn't manage to start the gluing today. 

  12301   Thu Jul 14 01:00:33 2016 KojiUpdateGeneralETMX ruby guided gluing

Today I took the picture of the glued ruby stand-off. The groove has not been invaded by the epoxy!

Attachment 1: ruby.jpg
ruby.jpg
  12300   Wed Jul 13 21:13:51 2016 KojiUpdateGeneralETMX guide rod gluing / ETMY Magnet gluing

[Eric, Koji]

  • ETMX: The Al guide rod has been glued on the mirror.
  • ETMY: The UR, LR, and SD magnets have been glued on the mirror.
  • For both, we are waiting for the glue getting cured.

Handing over message to the next step

  • ETMX: The arm of the fixture has probably been glued together with the guide rod. It has to be taken care when the fixture is removed.
    (See
    ETMX gluing section and Attachment 13)

ETMX: guide rod gluing (done) -> fixture unmounting side -> fixture setting -> magnet gluing -> suspend -> pitch balance -> ruby gluing -> air bake
ETMY: magnet (done) -> fixture unmounting -> air bake


ETMY gluing

Mirror transport:
- A transport setup was made with a donut holder for a 3" optic, glass jar, stain less tray, and a CS Stat zipbag. (Attachment 1)

Shimming:
- The magnets have been glued witht the gluing fixture. (Attachment 2)
- We checked the dimensions of the glued magnet and found that the thicker side has to be raised by 1mm. (We used the fact that the relative distance between the wire groove and the magnet is always the same.)
- The ETMs have 2.5deg wedge and this corresponds to 3.2mm height difference between the left and right edges. This meant that the thinner side had to be raised by 4.2mm.
- We used a 0.9mm Teflon sheet for the thicker side (Attachment 3) and two 2.2mm Teflon pieces for the thinner side (Attachment 4). For stabilization of the fixture, two Teflon tubes with a diameter of ~3mm are inserted to the top and bottom side of the mirror (Attachment 5).

Magnet orientation:
- Mirror orientation in the fixture (Attachment 6).
- It was confirmed that existing UR, LR, and Right SD magnets have the polarity of N facing out, S facing out, and N facing out. And we confirmed that this is consistent with ETMX and the procedure document (E970037)
- Along with the procedure document, we arranged the magnet-dumbbells UL, LL, and Left SD magnets to have S-out, N-out, and N-out. (Attachments 7, 8, and 9)

Gluing:
- In prior to gluing, all three dumbbells surfaces were cleaned by acetone and razor blade scrubbing.
- After the epoxy curing test (see below), the three magnet-dumbbell pairs have been glued on the mirror. A single dub of EP30-2 was applied to each dumbbell surface.

- Attachments 10, 11, and 12 shows how glue is spread at each joint.

ETMX gluing

Guide rod positioning:
- The longitudinal position of the guide rod was adjusted using the micrometer microscope such that it located at the center of the mirror thickness.
- The guide rod is not long enough to have the edges sticking out from the form of the fixture arm. Therefore only arm finger of the arm held the guide rod. 
- The height was adjusted to be 1.73mm (68mil) lower than the mirror scribe line. The mirror is fixed on the fixture upside down. So this bonds the guide rod above the scribe line.

Gluing:
- Then the epoxy was applied to the guide rod. The glue was applied to two edges of the rod, but capillary action spread the glue around the rod. It seemed that the fixture and the rod were connected with the glue. Care should be taken when the fixture is going to be removed. (Attachment 13)
- The top side (in the picture) where the stand-off will come is still relatively kept clean. So it must be OK for the stand off. If there is an issue, we can shave the epoxy with a razor blade.

Glue testing

- EP30-2 tends to fail to get cured. In order to check the mixture is properly made or not, we put a test piece into air bake oven.
- The procedure says, 200F 15min bake show if the glue is in a good shape or not.

- We have the temperature sensor setup on a air bake oven, but it seemed that the indicated temperature there is overestimate.
  The heating setting of 2 was enough to show the temp of 100degC although EP30-2 didn't get cured with this setting.

- Our experience says that heater setting of "5" makes the temperature ~90degC. On July 12nd, this setting showed the temp of 90degC. Today (July 13rd) it didn't. In the both cases, the epoxy got cured nicely. So we should use this setting.

Attachment 1: P7138742.jpg
P7138742.jpg
Attachment 2: gluing_fixture.jpg
gluing_fixture.jpg
Attachment 3: shimming3.jpg
shimming3.jpg
Attachment 4: shimming1.jpg
shimming1.jpg
Attachment 5: shimming2.jpg
shimming2.jpg
Attachment 6: gluing_fixture3.jpg
gluing_fixture3.jpg
Attachment 7: P7138747.jpg
P7138747.jpg
Attachment 8: P7138748.jpg
P7138748.jpg
Attachment 9: P7138750.jpg
P7138750.jpg
Attachment 10: P7138773.jpg
P7138773.jpg
Attachment 11: P7138774.jpg
P7138774.jpg
Attachment 12: P7138775.jpg
P7138775.jpg
Attachment 13: P7138772.jpg
P7138772.jpg
  12299   Wed Jul 13 15:35:56 2016 LydiaUpdateGeneralVent Progress - ETMY repositioned and removed

[Lydia, Johannes]

Took photos to document the original OSEM orientation and wrote down the serial numbers for each position. We removed the OSEMs, moved the suspension to the accessible side of the table and took out the optic, which was brought to the clean room to have the magnets reglued. The ETMY chamber is now closed up with the OSEMs and clamps inside on the table, and should not need to be reopened until the magnets have been reattached. 

  12298   Wed Jul 13 03:16:47 2016 KojiUpdateGeneralITMX dust

Multicolor flash light:

- It seems that the usb port charging doesn't work.

- There is a battery charger on Steve's desk. I set the batteries on it.

 

White LED flash light:

- I temporarily brought a compatible charger from WB. It's charging two batteries behind the LCD display on my desk.

 

ELOG V3.1.3-